One number is 12 less than another number. Their product is -32. What is the least value the smaller number could be?
Answer by oberobic(2304)

x*y = -32
x = y-12
substitute
(y-12)*y = -32
y^2 -12y = -32
y^2 -12y + 32 = 0
factor
(y-4)(y-8) = 0
so
y=4 or y=8
.
If y = 4, then x = 4-12 = -8
If y = 8, then x = 8-12 = -4
.
The question asks for the least value the smaller number could be: -8.